the sum of ages of kelvin and paul is 50 years in five years time the age of paul will be four times the age of kelvin now. how old is each boys

Answer by checkley79(3341)

K+P=50 OR K=50-P
(K+5)=4P
(50-P+5)=4P
55-P=4P
-4P-P=-55
-5P=-55
P=-55/-5
P=11 PAUL IS 10 YEARS OLD NOW.
K+11=50
K=50-11=39 KEVIN'S AGE NOW.
PROOF:
(39+5)=4*11
44=44